OpenStudy (anonymous):

The pressure in an engine cylinder increases from 7.6 atm to 13 atm as the gas is compressed from 0.50 L to 0.10 L. What is the expected cylinder temperature if the gas was initially 15°C?

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

@Somy

OpenStudy (somy):

oh dear im pretty sure @Abhisar u know it lol

OpenStudy (abhisar):

is it -174.47 C

OpenStudy (somy):

wait

OpenStudy (abhisar):

ohkay

OpenStudy (somy):

i love SI system so im gonna do it that way

OpenStudy (abhisar):

No that's wrong

OpenStudy (abhisar):

ohkay

OpenStudy (abhisar):

Its 985.26 K i think

OpenStudy (somy):

im getting approximately 569 Celsius

OpenStudy (somy):

oh wait i did it opposite way round lol

OpenStudy (somy):

ur first answer is right tho

OpenStudy (somy):

-174.47

OpenStudy (abhisar):

yes second one will be ryt

OpenStudy (abhisar):

Thankyou @Somy !

OpenStudy (somy):

\[ \frac{ P _{1} V _{1} }{ T _{1} } = \frac{ P _{2} V _{2} }{ T _{2} }\] \[\frac{ 7.6 \times 0.50 }{ 15+273 } =\frac{ 13\times 0.10 }{ T _{2} }\] \[T _{2}= \frac{ (13\times 0.10)\times(15+273) }{ 7.6\times0.50 }\] \[T _{2}= 98.5 K\] \[T _{2}= 98.5 K - 273= -174.5 °C \]

OpenStudy (somy):

u r welcome @Abhisar :*

OpenStudy (anonymous):

thank you!

OpenStudy (somy):

u r welcome :)